# Part II: Variational bootstrapping of the $\Lambda$CDM model

### v1

@@ Abstract
The method of variational bootstrapping, based on canonical variational completion, allows one to construct a Lagrangian for a physical theory depending on two sets of field variables, starting from a guess of the field equations for only one such set. This setup is particularly appealing for the construction of modified theories of gravity since one can take lessons from GR for the “educated guess” of the metric field equations; the field equations for the other fields are then fixed by the obtained Lagrangian (up to terms completely independent from the metric tensor).
In the present paper, we explore the applications of variational bootstrapping in the context of metric-affine theories of gravity. We find those metric-affine models which are, in a variational sense, closest to the $\Lambda$CDM model of cosmology. The method then allows to find “corrected” metric equations and to “bootstrap” the connection field equations.
The Lagrangians obtained via this method, though imposing some restricting criteria, encompass a wide variety of metric-affine models. In particular, they allow for quadratic metric-affine theories only if these avoid non-linear terms in the curvature tensor.
@@ I Introduction
Such generalizations of GR concern either the geometric structure of spacetime (i.e., the kinematical level), or the field equations (in other words, the dynamical level), or both.
@@
At the kinematical level, one of the simplest extension is metric-affine geometry, which allows for more freedom by relaxing the GR hypotheses on the spacetime connection (metricity, absence of torsion), [7, 8, 9]. In metric-affine gravity models, one can thus encode the dark sector into the “enhanced” geometry of spacetime [^1] Working directly at the dynamical level – without necessarily generalizing the spacetime structure – can be done by adding new fields in the theory such as, e.g., in scalar-tensor gravity [10]..
@@
At the dynamical level, one usually discusses modified theories of gravity by first postulating a Lagrangian, from which the field equations are derived. Even though this procedure is automatically justified as long as one expects a fundamental physical theory to admit a Lagrangian, ultimately what one typically wants to control are the properties of the field equations, rather than the ones of the Lagrangian. For this reason, one of us proposed in [11], together with D. Krupka, a systematic method, called canonical variational completion, allowing to go the other way around: starting with an approximate form (an “educated guess”) of the field equations, one can canonically construct a Lagrangian and subsequently, the “corrected” field equations. The obtained field equations are the closest ones to our original guess, in the sense that the correction terms to be added specifically measure the “obstruction from variationality” of this guess. The method of canonical variational completion was then upgraded in the first part of this work [12], to accommodate theories that involve more than one set of dynamical variables, e.g., a metric and a connection. We showed that from an “educated guess” of the field equations with respect to one variable (e.g., the metric) only, one can still canonically “bootstrap” a Lagrangian, up to boundary terms and to terms that are completely independent of the respective variable.
@@
In the present paper, we apply the method proposed in [12] to select those metric-affine Lagrangians producing the “closest to $\Lambda$CDM” metric equations. Our “educated guess” (or “seed”) for the metric equations is a system that formally resembles the Einstein equations with a cosmological constant (encoding dark energy, $w=-1$) and a term describing dark matter (with $w=0$). To obtain it, we first consider the form of general tensor fields fitting the dark energy or dark matter equation of state (EOS) under the assumption of cosmological symmetry. Next, we step back from cosmological symmetry and consider this expression as valid in full generality; this correspond to the idea of finding the most minimal extension of the $\Lambda$CDM equations. We then bootstrap this guess under the following hypotheses:
@@
  • The dark sector has a geometric (more precisely, metric-affine) origin. That is, the cosmological “constant” is actually a scalar function depending on the metric and on the connection components; similarly, the dark matter term can be completely expressed in terms of the metric and of the connection.
@@
  • Minimal coupling: In the seed equations, the dark sector terms are of order at most one.
@@
  • The metric tensor is used to build a volume form on spacetime and for raising/lowering indices of tensors.
  • Each of the dark sector terms can be expressed as a finite sum of homogeneous expressions (not necessarily of positive degrees) in the metric.
The second and third hypotheses above ensure on the one hand, that the obtained Lagrangians are generally covariant and, on the other hand, that all their non-boundary terms can be found from our “educated guess” of the metric equations only. This way, no input on the connection equations is necessary; these can be automatically “bootstrapped” from the obtained Lagrangian.
We find that a number of the existing metric-affine models fall into this class, e.g., those whose Lagrangian is the Ricci scalar of the independent connection and a subclass of quadratic metric-affine theories avoiding non-linear terms involving the curvature tensor of the independent connection.
The paper is structured as follows : In sections II and III, we fix our notations for the metric-affine geometry and briefly review some properties of the $\Lambda$CDM model that are necessary for our discussion. In section IV, we consider metric-affine theories of gravity based on modifications implemented directly at the level of GR field equations with aim of obtaining an “educated guess” for the (metric) field equation in a metric-affine extension of GR. Section V contains our main results. In section V.1, we remind the basics of the technique of variational bootstrapping presented in [12]. In section V.2, we explicit our hypothesis in the construction of our model before performing, in section V.3, the variational bootstrapping of the equations discussed in section IV – hence obtaining a Lagrangian for the full theory. Then, in section V.4, we study the field equations associated to the Lagrangian obtained in section V.3. In section V.5, we then comment on the obtained equations and on their link with previously formulated metric-affine models of gravity. Finally, in section VI, we present our conclusions and discuss further directions of research.
@@ II Geometric setup
In the following, we model spacetime by means of a triple $\left(M,\Gamma,\boldsymbol{g}\right)$, where $M$ is an arbitrary 4-dimensional differentiable manifold, $\Gamma$ is a generic linear connection on $M$ and $\boldsymbol{g}$ a metric with Lorentzian signature. Local coordinates on $M$ will be denoted by $x^{\mu}$ and the local basis vectors on $TM$, by $\partial_{\mu}$; when writing connection coefficients, the second lower index refers to the “derivative index” meaning that, for example, given a vector field $V=V^{\mu}\partial_{\mu}$ one has that $\nabla_{\mu}V^{\nu}\coloneqq\left(\nabla_{\partial_{\mu}}V\right)^{\nu}=\partial_{\mu}V^{\nu}+\Gamma^{\nu}_{\rho\mu}V^{\rho}$.
@@
Our metric-affine setup can equivalently (and more conveniently) be defined as the triple $\left(M,\mathbf{L},\boldsymbol{g}\right)$, where the distortion of $\Gamma$:
@@
is a tensor field of type $(1,2)$. The symbol $\circ$ designates the Levi-Civita connection associated to $\boldsymbol{g}$ and all its related quantities; for instance:
$\accentset{\circ}{\Gamma}_{\mu\nu}^{\rho}\coloneqq\frac{1}{2}g^{\rho\alpha}\left(\partial_{\nu}g_{\alpha\mu}+\partial_{\mu}g_{\nu\alpha}-\partial_{\alpha}g_{\mu\nu}\right),\,\,\Gamma_{\mu\nu}^{\rho}=\accentset{\circ}{\Gamma}_{\mu\nu}^{\rho}+L^{\rho}_{\phantom{\rho}\mu\nu}.$ (9)
In the following, for conciseness, we will express our relations in terms of the distortion tensor $\mathbf{L}$. If needed, all the relations can be translated in terms of the torsion $\mathbf{T}$ (respectively, the contortion $\mathbf{K}$) and non-metricity $\mathbf{Q}$ (respectively, disformation $\mathbf{D}$), using the relations below:
@@
$D^{\rho}_{\phantom{\rho}\mu\nu}\coloneqq-\frac{1}{2}g^{\rho\alpha}\left(Q_{\alpha\mu\nu}+Q_{\nu\alpha\mu}-Q_{\mu\nu\alpha}\right),$ (11)
@@
where, as usual, square brackets (resp. round brackets) placed around a group of indices denote antisymmetrization (resp. symmetrization) over those indices, with exception of indices placed between vertical bars.
@@ III Brief review of the $\Lambda$CDM model
This section briefly presents the restrictions imposed by cosmological symmetry upon tensors on the spacetime manifold and the geometric setup of the $\Lambda$CDM model; for more detail see [4, 5]. Even though most of the material of this section is not new, this allow us to clearly fix our notations and to raise an important – yet rarely emphasized – problem.
@@ III.1 Cosmologically symmetric tensor fields
Cosmological symmetry, understood as invariance under the six Killing vector fields defining it, imposes quite strong constraints on tensor fields of any rank, which we briefly present below; we direct the curious reader to Appendix A for more details and computations. In the following, we denote by $\left(x^{\mu}\right)=\left(t,r,\theta,\varphi\right)$, a set of local spherical coordinates over $M$ adapted to the symmetry.
@@ III.1.2 $(0,2)$-tensor fields
Similarly, a given $(0,2)$-tensor field $\boldsymbol{T}=T_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ will be cosmologically symmetric if and only if its components in the coordinate basis $\Set{\partial_{\mu}}$ satisfy
$\left(T_{\mu\nu}\right)=\operatorname{diag}\left(T_{tt}(t),\frac{T_{r}(t)}{1-kr^{2}},r^{2}T_{r}(t),r^{2}\sin(\theta)^{2}T_{r}(t)\right),$ (18)
where $k\in\Set{-1,0,1}$ is the sign of the curvature of the spatial slices. It is worth noting that, in deducing (18), one does not need to assume that $\boldsymbol{T}$ is symmetric; its antisymmetric part vanishes as a consequences of cosmological symmetry. From this relation, we naturally recover that the components of the most general cosmologically symmetric metric with Lorentz signature $\boldsymbol{g}=g_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ can be written as
@@ III.1.3 Differential 1-forms and $(1,1)$-type tensor fields
where we defined $\rho_{T}(t)\coloneqq T_{tt}(t)/N(t)^{2}$ and $P_{T}(t)\coloneqq T_{r}(t)/a(t)^{2}$.
@@ III.2 The $\Lambda$CDM model in a nutshell
where $w\in\mathbb{R}$ is a constant. For ordinary (dust) matter one has $w_{\text{dust}}=0$ (i.e. vanishing pressure) while for electromagnetic radiation (photons) $w_{\text{EM}}=1/3$ to ensure $T^{\text{(EM)}}{}^{\mu}_{\phantom{\mu}\mu}=0$. If we raise the first index of $\boldsymbol{T}^{\text{(m)}}$, we then get
@@
In the $\Lambda$CDM model, one considers the dark matter sector to be composed of “cold dark matter”. This cold dark matter represents, in this model, some exotic non-relativistic particles described by a perfect fluid energy-momentum tensor (23) for which $w_{\text{DM}}=0$. So
@@ III.3 The cosmological smokescreen
    Comparing (26) and (21) reveals that any type $(1,1)$-tensor field which obeys cosmological symmetry must look like a perfect fluid energy-momentum tensor.
These remarks, while supporting the parametrization of the energy-momentum tensor of ordinary matter via (23) subject to an EOS (25) with $w\in\left\{0,1/3\right\}$, can cast serious doubts on the interpretation of (28) and (29). If we further group together the “dark contributions” as
$T^{\text{(D)}}_{\mu\nu}\coloneqq T^{\text{(DE)}}_{\mu\nu}+T^{\text{(DM)}}_{\mu\nu}=-\Lambda/\kappa\ g_{\mu\nu}+\rho_{\text{DM}}\ u_{\mu}u_{\nu}$ (30)
and raise the first index, we recover the general expression (21) with associated “density” $\rho_{\text{D}}\coloneqq\rho_{\text{DM}}+\rho_{\text{DE}}=\rho_{\text{DM}}+\Lambda/\kappa$ and “pressure” $P_{\text{D}}\coloneqq P_{\text{DM}}+P_{\text{DE}}=-\Lambda/\kappa$.
A conservative interpretation of the $\Lambda$CDM model then seems to only indicate that there is a missing piece in Einstein equation $\accentset{\circ}{\mathbf{G}}=\kappa\boldsymbol{T}^{\text{(m)}}$, but does not allow to draw any stringent conclusion regarding the origin of the extra piece $\boldsymbol{T}^{\text{(D)}}$. It thus looks reasonable to consider the dark sector as a whole, as in (30), freed from the classical dark-matter+dark-energy interpretation.

### v2

@@ Abstract
The method of variational bootstrapping, based on canonical variational completion, allows one to construct a Lagrangian for a physical theory depending on two sets of field variables, starting from a guess of the field equations for only one such set. This setup is particularly appealing for the construction of modified theories of gravity, since one can take lessons from GR for an “educated guess” of the metric field equations; the field equations for the other fields are then fixed by the obtained Lagrangian (up to terms that are completely independent from the metric tensor).
In the present paper, we apply variational bootstrapping to determine metric-affine models which are, in a variational sense, closest to the $\Lambda$CDM model of cosmology. Starting from an “educated guess” that formally resembles the Einstein field equations with a cosmological “constant” (actually, a scalar function built from the metric and the connection) and a dark matter term, the method then allows to find “corrected” metric equations and to “bootstrap” the connection field equations.
Lagrangians obtained via this method, though imposing some restricting criteria, still encompass a wide variety of metric-affine models. In particular, they allow for a subclass of quadratic metric-affine theories restricted to linear terms in the curvature tensor.
@@ I Introduction
Such generalizations of GR concern the geometric structure of spacetime (i.e., the kinematical level), the field equations (in other words, the dynamical level), or both.
@@
At the kinematical level, one of the simplest extensions is metric-affine geometry, which allows for more freedom by relaxing the GR hypotheses on the spacetime connection (metricity, absence of torsion), [7, 8, 9]. In metric-affine gravity models, one can thus encode the dark sector into the “enhanced” geometry of spacetime [^1] Working directly at the dynamical level – without necessarily generalizing the spacetime structure – can be done by adding new fields in the theory such as, e.g., in scalar-tensor gravity [10]..
@@
At the dynamical level, one usually discusses modified theories of gravity by first postulating a Lagrangian, from which the field equations are then derived. Even though this procedure is automatically justified as long as one expects a fundamental physical theory to admit a Lagrangian, ultimately what one typically wants to control are the properties of the field equations, rather than the ones of the Lagrangian. For this reason, one of us proposed in [11], together with D. Krupka, a systematic method, called canonical variational completion, allowing to go the other way around: starting with an approximate form (an “educated guess”) of the field equations, one can canonically construct a Lagrangian and subsequently, the “corrected” field equations. The obtained field equations are the closest ones to our original guess, in the sense that the correction terms to be added specifically measure the “obstruction from variationality” of this guess. The method of canonical variational completion was then upgraded in the first part of this work [12], to accommodate theories that involve more than one set of dynamical variables, e.g., a metric and a connection. We showed that from an “educated guess” of the field equations with respect to one variable (e.g., the metric) only, one can still canonically “bootstrap” a Lagrangian, up to boundary terms and to terms that are completely independent of the respective variable.
@@
In the present paper, we apply the method proposed in [12] to select those metric-affine Lagrangians producing the “closest to $\Lambda$CDM” metric equations. Our “educated guess” (or “seed”) for the metric equations is a system that formally resembles the Einstein equations with a cosmological constant (encoding dark energy, $w=-1$) and a term describing dark matter (with $w=0$), constructed following several ideas which we list below:
  1. 1.
    General covariance. The seed equations must be generally covariant.
  2. 2.
    The metric-affine hypothesis. The dark sector has a geometric (more precisely, metric-affine) origin.
@@
    The above, together with general covariance, imply that the cosmological “constant” is actually a scalar function depending on the metric and on the connection components; similarly, the dark matter term is a tensor field which can be completely expressed in terms of the metric and of the connection.
  3. 3.
    Role of the metric. The volume form on spacetime and raising/lowering indices of tensors are the standard ones in terms of the spacetime metric tensor.
  4. 4.
    Minimality of the extension. By minimality, we will actually mean three different aspects:
@@
    • Minimal order: In the seed equations, the dark sector terms are of order at most one.
@@
    • No ad-hoc terms: To ensure this, we first consider the particular situation of cosmological symmetry – and in this situation, we determine the most general form of (0,2)-type tensor fields fitting the dark energy, respectively, dark matter equations of state (EOS) [^2] Actually, we point out a much stronger statement: any cosmologically symmetric (0,2)-type tensor field can be split into a sum of two terms which formally obey the equations of state of dark energy, respectively, of dark matter.. Next, stepping back from cosmological symmetry, we consider the sum of the two obtained tensor fields (without adding any other, “ad hoc”, terms) as our “educated guess”, for the fully general case.
    • The finite sum ansatz: Each of the dark sector terms can be expressed as a finite sum of homogeneous expressions (not necessarily of positive degrees) in the metric and its partial derivatives.
      To make sense of the last hypothesis, we must bear in mind that the metric is used in the dark sector terms in three ways: to construct connection coefficients/curvature components, to raise indices, respectively, to construct the volume form. Each of these operations employs a finite number of metric tensor components or partial derivatives thereof.
The above hypotheses ensure, on the one hand, that the obtained Lagrangians are generally covariant and, on the other hand, that all their non-boundary terms can be found from our “educated guess” (or seed) of the metric equations only. This way, no input on the connection equations is necessary; these can be automatically “bootstrapped” from the obtained Lagrangian.
We find that a number of the existing metric-affine models do satisfy the above minimality requests, e.g., those whose Lagrangian is the Ricci scalar of the independent connection and a subclass of quadratic metric-affine theories avoiding non-linear terms in the curvature tensor of the connection.
The paper is structured as follows: In sections II and III, we fix our notations for the metric-affine geometry and briefly review some properties of the $\Lambda$CDM model that are necessary for our discussion. In section IV, we consider metric-affine theories of gravity based on modifications implemented directly at the level of GR field equations, with the aim of obtaining an “educated guess” for the metric field equation in a metric-affine extension of GR. Section V contains our main results, as follows. In subsection V.1, we remind the basics of the technique of variational bootstrapping presented in [12]. Then, in subsection V.2, we explain our hypotheses in the construction of our “educated guess”, before performing, in subsection V.3, its variational bootstrapping, leading to a Lagrangian (actually, a class of Lagrangians) for the full theory. Then, in subsection V.4, we determine the field equations for the obtained Lagrangians. In subsection V.5, we then comment on the obtained equations and on their link with previously formulated metric-affine models of gravity. Finally, in section VI, we present our conclusions and discuss further directions of research.
@@ II Geometric setup
In the following, we model spacetime by means of a triple $\left(M,\Gamma,\boldsymbol{g}\right)$, where $M$ is an arbitrary 4-dimensional differentiable manifold, $\Gamma$ is a generic linear connection on $M$ and $\boldsymbol{g}$ a metric with Lorentzian signature. Local coordinates on $M$ will be denoted by $x^{\mu}$ and the natural (coordinate) local basis vectors on $TM$, by $\partial_{\mu}$. When writing connection coefficients, the second lower index refers to the “derivative index” meaning that, for example, given a vector field $V=V^{\mu}\partial_{\mu}$ one has
$\nabla_{\mu}V^{\nu}\coloneqq\left(\nabla_{\partial_{\mu}}V\right)^{\nu}=\partial_{\mu}V^{\nu}+\Gamma^{\nu}_{\rho\mu}V^{\rho}.$
@@
Our metric-affine setup can equivalently (and more conveniently, for our purposes) be defined as the triple $\left(M,\mathbf{L},\boldsymbol{g}\right)$, where the distortion of $\Gamma$:
@@
is a tensor field of type $(1,2)$, with components $L^{\rho}_{\phantom{\rho}\mu\nu}$ in a coordinate basis. The symbol $\circ$ designates the Levi-Civita connection associated to $\boldsymbol{g}$ and all its related quantities; for instance:
$\accentset{\circ}{\Gamma}_{\mu\nu}^{\rho}\coloneqq\frac{1}{2}g^{\rho\alpha}\left(\partial_{\nu}g_{\mu\alpha}+\partial_{\mu}g_{\alpha\nu}-\partial_{\alpha}g_{\mu\nu}\right),\,\,\Gamma_{\mu\nu}^{\rho}=\accentset{\circ}{\Gamma}_{\mu\nu}^{\rho}+L^{\rho}_{\phantom{\rho}\mu\nu}.$ (9)
In the following, for conciseness, we will express our relations in terms of the distortion tensor $\mathbf{L}$. If needed, all the relations can be translated in terms of the torsion $\mathbf{T}$ (respectively, the contortion $\mathbf{K}$) and non-metricity $\mathbf{Q}$ (respectively, disformation $\mathbf{D}$) using the relations below, [9]:
@@
$D^{\rho}_{\phantom{\rho}\mu\nu}\coloneqq-\frac{1}{2}g^{\rho\alpha}\left(Q_{\nu\alpha\mu}-Q_{\mu\nu\alpha}+Q_{\alpha\mu\nu}\right),$ (11)
@@
where, as usual, square brackets (resp. round brackets) placed around a group of indices denote total antisymmetrization (resp. symmetrization) over those indices, with exception of indices placed between vertical bars.
@@ III Brief review of the $\Lambda$CDM model
This section briefly presents the restrictions imposed by cosmological symmetry upon tensors on the spacetime manifold and the geometric setup of the $\Lambda$CDM model; for more detail see [4, 5]. Even though most of the material of this section is not new, this allows us to clearly fix our notations and to recall an important – yet rarely emphasized – problem.
@@ III.1 Cosmologically symmetric tensor fields
Cosmological symmetry, understood as invariance under the six Killing vector fields defining it, imposes quite strong constraints on tensor fields of any rank, which we briefly present below; these six symmetry generators are explicitly displayed in Appendix A.2. In the following, we denote by $\left(x^{\mu}\right)=\left(t,r,\theta,\varphi\right)$, a set of local spherical coordinates over $M$ adapted to the symmetry.
@@ III.1.2 $(0,2)$-tensor fields
Similarly, a given $(0,2)$-tensor field $\boldsymbol{T}=T_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ is cosmologically symmetric if and only if its components in the coordinate basis $\Set{\partial_{\mu}}$ satisfy
$\left(T_{\mu\nu}\right)=\operatorname{diag}\left(T_{t}(t),\frac{T_{r}(t)}{1-kr^{2}},r^{2}T_{r}(t),r^{2}\sin(\theta)^{2}T_{r}(t)\right),$ (18)
where $k\in\Set{-1,0,1}$ is the sign of the curvature of the spatial slices. It is worth noting that, in deducing (18), one does not need to assume that $\boldsymbol{T}$ is symmetric; its antisymmetric part vanishes as a consequence of cosmological symmetry. From this relation, we naturally recover that the components of the most general cosmologically symmetric metric with Lorentz signature $\boldsymbol{g}=g_{\mu\nu}\mathrm{d}x^{\mu}\otimes\mathrm{d}x^{\nu}$ can be written as
@@
Of course, in the above, the “lapse function” $N(t)$ can be absorbed into the definition of the cosmological time coordinate $t$, so that $g_{00}=-1$. We preferred, yet to leave it general for the sake of flexibility (e.g., if one wants to work in terms of the conformal time $\eta$).
@@ III.1.3 Differential 1-forms and $(1,1)$-type tensor fields
where we defined $\rho_{T}(t)\coloneqq T_{t}(t)/N(t)^{2}$ and $P_{T}(t)\coloneqq T_{r}(t)/a(t)^{2}$.
@@ III.2 The $\Lambda$CDM model in a nutshell
where $w\in\mathbb{R}$ is a constant. For instance, for ordinary (dust) matter one has $w_{\text{dust}}=0$ (that is, vanishing pressure), while for electromagnetic radiation (photons), the equation of state is $w_{\text{EM}}=1/3$ (ensuring that $T^{\text{(EM)}}{}^{\mu}_{\phantom{\mu}\mu}=0$).
If we raise the first index of $\boldsymbol{T}^{\text{(m)}}$, we then get
@@
In the $\Lambda$CDM model, one considers the dark matter sector to be composed of cold dark matter, made of some hypothesized non-relativistic particles, described by a perfect fluid energy-momentum tensor (23) for which $w_{\text{DM}}=0$. So
@@ III.3 The cosmological smokescreen
    Comparing (26) and (21) reveals that any type $(1,1)$-tensor field which obeys cosmological symmetry must formally look like a perfect fluid energy-momentum tensor.
Obviously, ordinary matter obeys the above criteria. Assuming cosmological symmetry, this should also be the case for the dark sector energy-momentum tensor (of whatever origin). It then turns out that any cosmologically symmetric extra piece $T^{\text{(D)}}_{\mu\nu}$ of the energy-momentum tensor can be formally split into a “dark energy” and a “dark matter” part, as seen below:
$T^{\text{(D)}}_{\mu\nu}\coloneqq T^{\text{(DE)}}_{\mu\nu}+T^{\text{(DM)}}_{\mu\nu}=-\Lambda/\kappa\ g_{\mu\nu}+\rho_{\text{DM}}\ u_{\mu}u_{\nu},$ (30)
where $\Lambda$ is, this time, no longer a constant, but a scalar function of $t$. Indeed, raising the first index, we recover the general expression (21) with associated “density” $\rho_{\text{D}}\coloneqq\rho_{\text{DM}}+\rho_{\text{DE}}=\rho_{\text{DM}}+\Lambda/\kappa$ and “pressure” $P_{\text{D}}\coloneqq P_{\text{DM}}+P_{\text{DE}}=-\Lambda/\kappa$. A conservative interpretation of the $\Lambda$CDM model thus only indicates that there is a missing piece in Einstein equation $\accentset{\circ}{\mathbf{G}}=\kappa\boldsymbol{T}^{\text{(m)}}$, but does not allow to draw any stringent conclusion regarding the physical interpretation of the extra piece $\boldsymbol{T}^{\text{(D)}}$. It then looks reasonable to consider the dark sector in full generality, as in (30). The important fact, for us, is that the dark sector energy momentum tensor (of whatever origin) has, in cosmological symmetry, the general form (30) and one has two degrees of freedom in its choice, represented by the scalars $\Lambda$ and $\rho_{DM}$.